# Divergences induced by dual subtractive and divisive normalizations of exponential families and their convex deformations

### 20 Dec 2023 v1

@@ Abstract
Exponential families are statistical models which are the workhorses in statistics, information theory, and machine learning. An exponential family can either be normalized subtractively by its cumulant function or equivalently normalized divisively by its partition function. Both subtractive and divisive normalizers are strictly convex and smooth functions inducing pairs of Bregman and Jensen divergences. It is well-known that skewed Bhattacharryya distances between probability densities of an exponential family amounts to skewed Jensen divergences induced by the cumulant function between their corresponding natural parameters, and in limit cases that the sided Kullback-Leibler divergences amount to reverse-sided Bregman divergences. In this note, we first show that the $\alpha$-divergences between unnormalized densities of an exponential family amounts scaled $\alpha$-skewed Jensen divergences induced by the partition function. We then show how comparative convexity with respect to a pair of quasi-arithmetic means allows to deform convex functions and define dually flat spaces with corresponding divergences when ordinary convexity is preserved.
Keywords: convex duality; exponential family; Bregman divergence; Jensen divergence; comparative convexity; quasi-arithmetic means; information geometry.
@@ 1 Introduction
In information geometry [3], any strictly convex and smooth function induces a dually flat space (DFS) with a canonical divergence which can be expressed in charts either as dual Bregman divergences [6] or equivalently as dual Fenchel-Young divergences [23]. For example, the cumulant function of an exponential family [7] generates a DFS: An exponential family manifold [29, 19] with the canonical divergence yielding the reverse Kullback-Leibler divergence [18]. Another typical example is the negentropy of a mixture family inducing a DFS: A mixture family manifold with the canonical divergence yielding the Kullback-Leibler divergence [23]. Any strictly convex and smooth function also induces a family of scaled skewed Jensen divergences [32, 21] which includes in limit cases the sided forward and reverse Bregman divergences.
In §2, we show that they are two equivalent approaches to normalize an exponential family: Either by its cumulant function or by its partition function. Since both cumulant and partition functions are convex, they induce (i) families of scaled skewed Jensen divergences and (ii) dually flat spaces with corresponding Bregman divergences corresponding to statistical divergences (Proposition 4). In §3, we recall the well-known result that the statistical $\alpha$-skewed Bhattacharyya distances between probability densities of an exponential family amount to a scaled $\alpha$-skewed Jensen divergence between their natural parameters. In §4, we prove that the $\alpha$-divergences between unnormalized densities of a exponential family amount to scaled $\alpha$-skewed Jensen divergence between their natural parameters (Proposition 5). More generally, we explain in section 5 how to deform a convex function using comparative convexity [17]: When ordinary convexity of the deformed convex function is preserved, we thus get new skew Jensen divergences and DFSs with corresponding Bregman divergences. Section 6 concludes this work with a discussion.
@@ 2 Dual subtractive and divisive normalizations of exponential families
An exponential family [7, 22] is a set of probability distributions $\mathcal{P}=\{P_{\lambda}\ :\ \lambda\in\Lambda\}$ all dominated by $\mu$ such that their Radon-Nikodym densities $p_{\lambda}(x)=\frac{\mathrm{d}P_{\lambda}}{\mathrm{d}\mu}(x)$ can be expressed as
@@
Exponential families include many well-known distributions: For example, the Bernoulli distributions, the Gaussian or normal distributions, the Gamma and Beta distributions, the Wishart distributions, the Poisson distributions, the Rayleigh distributions, etc. Furthermore, any two probability measures $Q$ and $R$ with densities $q$ and $r$ with respect to dominating measure $\mu=\frac{Q+R}{2}$ define an exponential family :
@@
called the likelihood ratio exponential family [13] because the sufficient statistic is $t(x)=\log\frac{q(x)}{r(x)}$ (with auxiliary carrier term $h(x)=r(x)$) or the Bhattacharryya arc since the cumulant function of $\mathcal{P}_{Q,R}$ is expressed as the negative of the skewed Bhattacharryya distances [16, 21].
@@
In the context of $\lambda$-deformed exponential families [33] which generalize exponential families, function $Z(\theta)$ is called the divisive normalization factor (Eq. 1) and function $F(\theta)$ is called the subtractive normalization factor (Eq. 2).
The order $m$ of an exponential family is the dimension of the sufficient statistic $t(x)$. The full natural parameter space $\Theta$ is defined by $\Theta=\{\theta\ :\ \int{\tilde{p}}_{\theta}(x)\mathrm{d}\mu(x)<\infty\}$. The exponential family is said full when $\theta$ is allowed to range in the full parameter space $\Theta$ (curved exponential families [3] have natural parameters $c(\theta)$ ranging in a subset of $\Theta$), and regular when $\Theta$ is an open (convex) domain. A full regular exponential family such that both $t(x)=x$ and $k(x)=0$ is called a natural exponential family. Without loss of generality, we consider natural exponential families or exponential families with $k(x)=0$ in the remainder.
We shall prove that both the cumulant function $F(\theta)$ and the partition function $Z(\theta)$ are strictly convex functions [7].
Proposition 1
@@
Proof:
Let $\Theta$ denote the natural parameter space:
since both $\int\exp({\left\langle\theta_{0},x\right\rangle})\mathrm{d}\mu(x)<\infty$ and $\int\exp({\left\langle\theta_{1},x\right\rangle})\mathrm{d}\mu(x)<\infty$ because $\theta_{0}$ and $\theta_{1}$ both belong to $\Theta$. Hence, we have shown that $\Theta$ is convex. $\square$
Proposition 2
@@
Proof:
Eq. 4 can be rewritten as
Now, a function $Z(\theta)$ is strictly log-convex iff $\log Z(\theta)$ is strictly convex [17]. Since $Z(\theta)=\exp(F(\theta))$, we deduce that $Z(\theta)$ is positive and strictly log-convex since $\log Z(\theta)=F(\theta)$ is strictly convex. $\square$ It can be shown that the cumulant and partition functions are smooth $C^{\infty}$ analytic functions [7].
A remarkable property is that log-convex functions are also convex:
Proposition 3
A strictly log-convex $C^{2}$ function $Z:\Theta\subset\mathbb{R}^{m}\rightarrow\mathbb{R}$ is strictly convex.
Proof:
  • In the univariate case $m=1$, consider the log-convex function $Z(\theta)=\exp({F(\theta)})$ for $F(\theta)$ a strictly convex function with $F^{\prime\prime}(\theta)>0$. Then we have $Z^{\prime\prime}(\theta)=F^{\prime\prime}(\theta)e^{F(\theta)}+(F^{\prime}(\theta))^{2}e^{F(\theta)}>0$ since $F(\theta)$ is strictly convex.
  • In the multivariate case $m>1$, $C^{2}$ function $F(\theta)$ is strictly convex iff its Hessian is positive-definite. Let $Z(\theta)$ be $C^{2}$ strictly log-convex, i.e., $F(\theta)=\log Z(\theta)$ is strictly convex. Since $Z(\theta)=e^{F(\theta)}$ for $\nabla^{2}F(\theta)\succ 0$, we need to check that $\nabla^{2}Z(\theta)\succ 0$. Let $\partial_{l}=\frac{\partial}{\partial\theta_{l}}$. Then we have
    or in matrix form
    Since (i) $e^{F(\theta)}>0$, (ii) $\nabla^{2}F(\theta)\succ 0$ and (iii) $\nabla F(\theta)(\nabla F(\theta))^{\top}\succeq 0$ (since a matrix outer-product is positive semi-definite: $\theta^{\top}\nabla F(\theta)(\nabla F(\theta))^{\top}\theta=(\theta^{\top}\nabla F(\theta))^{2}\geq 0$), we deduce that $\nabla^{2}Z(\theta)$ is positive-definite. Hence, $Z(\theta)$ is strictly convex.
$\square$
@@
The converse is not necessarily true: Some convex functions are not log-convex, and thus the class of strictly log-convex functions is a proper subclass of strictly convex functions. For example, $\theta^{2}$ is convex but log-concave since $(\log\theta^{2})^{\prime\prime}=-\frac{2}{\theta^{2}}<0$ (Figure 1).
Proposition 4
The cumulant and partition functions of an exponential family are strictly convex and smooth functions inducing a pair of dually flat spaces with corresponding Bregman divergences, and a pair of families of skewed Jensen divergences.
Proof:
A log-convex function $Z$ induces a pair $(B_{Z},B_{\log Z}=B_{F})$ of Bregman divergences [6]:
@@
and a pair of $\alpha$-skewed Jensen divergences [8] for $\alpha\in(0,1)$:
@@
$\square$
For a strictly convex function $F(\theta)$, we define the symmetric Jensen divergence $J_{F}(\theta_{1},\theta_{2})=J_{F,\frac{1}{2}}(\theta_{1},\theta_{2})$.
Notice that function $F(\theta)$ is called the cumulant function because when $X\sim p_{\theta}(x)$ is a random variable following a probability distribution of an exponential family, function $F(\theta)$ appears in the cumulant generating function of $X$: $K_{X}(t)=\log E_{X}[e^{{\left\langle t,X\right\rangle}}]=F(\theta+t)-F(\theta)$. The cumulant function is also called the log-normalizer or log-partition function in statistical physics.
Remark 1
@@
$\Leftrightarrow$ $F(\theta_{1})\geq F(\theta_{2})+{\left\langle\theta_{1}-\theta_{2},\nabla F(\theta)\right\rangle},$
@@
$\Leftrightarrow$ $\log Z(\theta_{1})\geq\log Z(\theta_{2})+{\left\langle\theta_{1}-\theta_{2},\frac{\nabla Z(\theta_{2})}{Z(\theta_{2})}\right\rangle}.$
@@
$J_{F,\alpha}^{s}(\theta_{1}:\theta_{2})=\left\{\begin{array}[]{ll}\frac{1}{\alpha(1-\alpha)}J_{F,\alpha}(\theta_{1}:\theta_{2}),&\alpha\in\mathbb{R}\backslash\{0,1\},\\ B_{F}(\theta_{1}:\theta_{2}),&\alpha=0,\\ B_{F}^{*}(\theta_{1}:\theta_{2})=B_{F}(\theta_{2}:\theta_{1}),&\alpha=1.\end{array}\right.,$ (10)
@@
Next, we first recall the connections between these parameter Jensen and Bregman divergences and statistical divergence counterparts between probability densities in §3. We will then introduce the novel connections between these parameter divergences and $\alpha$-divergences between unnormalized densities in §4.
@@ 3 Divergences related to the cumulant function
$D_{B,\alpha}^{s}(p:q)=-\frac{1}{\alpha(1-\alpha)}\log\int p^{\alpha}q^{1-\alpha}\mathrm{d}\mu.$
When both probability densities belong to the same exponential family =$\mathcal{E}=\{p_{\theta}(x)\ :\ \theta\in\Theta\}$ with cumulant $F(\theta)$, we have:
@@
The proof follows by first considering the $\alpha$-skewed Bhattacharyya similarity coefficient $\rho_{\alpha}(p,q)=\int p^{\alpha}q^{1-\alpha}\mathrm{d}\mu$. We have $\rho_{\alpha}(p_{\theta_{1}}:p_{\theta_{2}})=\exp(-J_{F,\alpha}(\theta_{1}:\theta_{2}))$, see [21].
Proof:
$\rho_{\alpha}(p_{\theta_{1}}:p_{\theta_{2}})$ $=$ $\int\exp({\left\langle\theta_{1},x\right\rangle}-F(\theta_{1}))^{\alpha}\,\exp({\left\langle\theta_{2},x\right\rangle}-F(\theta_{2}))^{1-\alpha}\mathrm{d}\mu,$
$=$ $\int\exp({\left\langle\alpha\theta_{1}+(1-\alpha)\theta_{2}),x\right\rangle})\,\exp(-(\alpha F(\theta_{1})+(1-\alpha)F(\theta_{2}))\mathrm{d}\mu.$
@@
Since $\bar{\theta}\in\Theta$, we have $\int\exp({\left\langle\bar{\theta},x\right\rangle}-F(\bar{\theta}))\mathrm{d}\mu=1$, and therefore $\rho_{\alpha}(p_{\theta_{1}}:p_{\theta_{2}})=\exp(-J_{F,\alpha}(\theta_{1}:\theta_{2}))$. $\square$
@@
This formula has been further generalized in [20] by considering truncations of exponential family densities: Let $\mathcal{X}_{1}\subseteq\mathcal{X}_{2}\subseteq\mathcal{X}$ and $\mathcal{E}_{1}=\{1_{\mathcal{X}_{1}}(x)\,p_{\theta}(x)\}$, $\mathcal{E}_{2}=\{1_{\mathcal{X}_{2}}(x)\,q_{\theta^{\prime}}(x)\}$ be two truncated families of $\mathcal{X}$ with cumulant functions $F_{1}(\theta)=\log\int_{\mathcal{X}_{1}}\exp({\left\langle t(x),\theta\right\rangle})\mathrm{d}\mu$ and $F_{2}(\theta^{\prime})=\log\int_{\mathcal{X}_{2}}\exp({\left\langle t(x),\theta\right\rangle})\mathrm{d}\mu\geq F_{1}(\theta^{\prime})$. Then we have
@@
Truncated exponential families are exponential families which may not be regular [9] (i.e., parameter space $\Theta$ may not be open).

### 18 Jan 2024 v2

@@ Abstract
Exponential families are statistical models which are the workhorses in statistics, information theory, and machine learning among others. An exponential family can either be normalized subtractively by its cumulant or free energy function or equivalently normalized divisively by its partition function. Both subtractive and divisive normalizers are strictly convex and smooth functions inducing pairs of Bregman and Jensen divergences. It is well-known that skewed Bhattacharryya distances between probability densities of an exponential family amounts to skewed Jensen divergences induced by the cumulant function between their corresponding natural parameters, and in limit cases that the sided Kullback-Leibler divergences amount to reverse-sided Bregman divergences. In this paper, we first show that the $\alpha$-divergences between unnormalized densities of an exponential family amounts to scaled $\alpha$-skewed Jensen divergences induced by the partition function. We then show how comparative convexity with respect to a pair of quasi-arithmetic means allows to deform both convex functions and their arguments, and thereby define dually flat spaces with corresponding divergences when ordinary convexity is preserved.
Keywords: convex duality; exponential family; Bregman divergence; Jensen divergence; Bhattacharyya distance; Rényi divergence; $\alpha$-divergences; comparative convexity; log convexity; exponential convexity; quasi-arithmetic means; information geometry.
@@ 1 Introduction
In information geometry [4], any strictly convex and smooth function induces a dually flat space (DFS) with a canonical divergence which can be expressed in charts either as dual Bregman divergences [9] or equivalently as dual Fenchel-Young divergences [33]. For example, the cumulant function of an exponential family [11] (also called free energy) generates a DFS: An exponential family manifold [36] with the canonical divergence yielding the reverse Kullback-Leibler divergence. Another typical example is the negative entropy (negentropy for short) of a mixture family inducing a DFS: A mixture family manifold with the canonical divergence yielding the Kullback-Leibler divergence [33]. Any strictly convex and smooth function also induces a family of scaled skewed Jensen divergences [42, 32] which includes in limit cases the sided forward and reverse Bregman divergences.
In §2, we show that there are two equivalent approaches to normalize an exponential family: Either by its cumulant function or by its partition function. Since both cumulant and partition functions are convex, they induce (i) families of scaled skewed Jensen divergences and (ii) dually flat spaces with corresponding Bregman divergences corresponding to statistical divergences (Proposition 1). In §3, we recall the well-known result that the statistical $\alpha$-skewed Bhattacharyya distances between probability densities of an exponential family amount to a scaled $\alpha$-skewed Jensen divergence between their natural parameters. In §4, we prove that the $\alpha$-divergences between unnormalized densities of a exponential family amount to scaled $\alpha$-skewed Jensen divergence between their natural parameters (Proposition 5). More generally, we explain in section 5 how to deform a convex function using comparative convexity [29]: When ordinary convexity of the deformed convex function is preserved, we thus get new skew Jensen divergences and DFSs with corresponding Bregman divergences. Section 6 concludes this work with a discussion.
@@ 2 Dual subtractive and divisive normalizations of exponential families
An exponential family [11] is a set of probability distributions $\mathcal{P}=\{P_{\lambda}\ :\ \lambda\in\Lambda\}$ all dominated by $\mu$ such that their Radon-Nikodym densities $p_{\lambda}(x)=\frac{\mathrm{d}P_{\lambda}}{\mathrm{d}\mu}(x)$ can be expressed as
@@
Exponential families include many well-known distributions: For example, the Bernoulli distributions, the Gaussian or normal distributions, the Gamma and Beta distributions, the Wishart distributions, the Poisson distributions, the Rayleigh distributions, etc. The categorical distributions (i.e., discrete distributions on a finite alphabet sample space) form an exponential family too, and exponential families can universally model arbitrarily closely any smooth density [12, 30]. Furthermore, any two probability measures $Q$ and $R$ with densities $q$ and $r$ with respect to dominating measure $\mu=\frac{Q+R}{2}$ define an exponential family :
@@
called the likelihood ratio exponential family [19] because the sufficient statistic is $t(x)=\log\frac{q(x)}{r(x)}$ (with auxiliary carrier term $h(x)=r(x)$) or the Bhattacharryya arc since the cumulant function of $\mathcal{P}_{Q,R}$ is expressed as the negative of the skewed Bhattacharryya distances [24, 32]. In machine learning, undirected graphical models [39] and energy-based models [28] including Markov random fields [25] and conditional random fields are exponential families [13].
@@
We may view Eq. 1 as an exponential tilting [15] of density $h(x)\mathrm{d}\mu(x)$.
In the context of $\lambda$-deformed exponential families [43] which generalize exponential families, function $Z(\theta)$ is called the divisive normalization factor (Eq. 1) and function $F(\theta)$ is called the subtractive normalization factor (Eq. 2). Notice that function $F(\theta)$ is called the cumulant function because when $X\sim p_{\theta}(x)$ is a random variable following a probability distribution of an exponential family, function $F(\theta)$ appears in the cumulant generating function of $X$: $K_{X}(t)=\log E_{X}[e^{{\left\langle t,X\right\rangle}}]=F(\theta+t)-F(\theta)$. The cumulant function is also called the log-normalizer or log-partition function in statistical physics. Since $Z>0$ and $F=\log Z$, we deduce that $F\geq Z$ since $\log x\leq x$ for $x>0$.
The order $m$ of an exponential family is the dimension of the sufficient statistic $t(x)$. The full natural parameter space $\Theta\in\mathbb{R}^{m}$ is defined by $\Theta=\{\theta\ :\ \int{\tilde{p}}_{\theta}(x)\mathrm{d}\mu(x)<\infty\}$. The exponential family is said full when $\theta$ is allowed to range in the full parameter space $\Theta$ (curved exponential families [4] have natural parameters $c(\theta)$ ranging in a subset of $\Theta$), and regular when $\Theta$ is an open (convex) domain. A full regular exponential family such that both $t(x)=x$ and $k(x)=0$ is called a natural exponential family. Without loss of generality, we consider natural exponential families or exponential families with $k(x)=0$ in the remainder.
It is well-known that the cumulant function $F(\theta)$ is a strictly convex function and that the partition function $Z(\theta)$ is strictly log-convex [7] (see Appendix A):
Proposition 1 ([7]).
@@
Proposition 2 ([7]).
@@
It can be shown that the cumulant and partition functions are smooth $C^{\infty}$ analytic functions [11].
A remarkable property is that strictly log-convex functions are also strictly convex:
Proposition 3.
A strictly log-convex function $Z:\Theta\subset\mathbb{R}^{m}\rightarrow\mathbb{R}$ is strictly convex.
Proof.
By definition, function $Z(\theta)$ is strictly log-convex if and only if:
$\forall\theta_{0}\not=\theta_{1},\quad Z(\alpha\theta_{0}+(1-\alpha\theta_{1}))<Z(\theta_{0})^{\alpha}\,Z(\theta_{1})^{1-\alpha},$ (3)
i.e., by taking the logarithm on both sides of the inequality, $F=\log Z$ is strictly convex:
$\forall\theta_{0}\not=\theta_{1},\quad\log Z(\alpha\theta_{0}+(1-\alpha)\theta_{1})$ $<$ $\alpha\log Z(\theta_{0})+(1-\alpha)\log Z(\theta_{1}),$
$\Leftrightarrow\quad F(\alpha\theta_{0}+(1-\alpha)\theta_{1})$ $<$ $\alpha F(\theta_{0})+(1-\alpha)F(\theta_{1}).$
Since $f(x)=\exp(x)$ is strictly convex (because $f^{\prime\prime}(x)=\exp(x)>0$), we have for all $\alpha\in(0,1)$:
$f(\alpha F(\theta_{0})+(1-\alpha)F(\theta_{1}))<\alpha f(F(\theta_{0}))+(1-\alpha)f(F(\theta_{1})).$
Letting $F(\theta)=\log Z(\theta)$ in the above inequality, we get:
$\exp(\alpha\log Z(\theta_{0})+(1-\alpha)\log Z(\theta_{1})$ $<$ $\alpha\exp(\log Z(\theta_{0}))+(1-\alpha)\exp(\log Z(\theta_{1})),$ (4)
$Z(\theta_{0})^{\alpha}\,Z(\theta_{1})^{1-\alpha}$ $<$ $\alpha Z(\theta_{0})+(1-\alpha)Z(\theta_{1}),$ (5)
and therefore we get from Eq. 3 and Eq. 5:
$\forall\theta_{0}\not=\theta_{1},Z(\alpha\theta_{0}+(1-\alpha\theta_{1}))<Z(\theta_{0})^{\alpha}\,Z(\theta_{1})^{1-\alpha}<\alpha Z(\theta_{0})+(1-\alpha)Z(\theta_{1}).$ (6)
That is, $Z$ is strictly convex.
An alternative proof may use the weighted arithmetic-mean geometric mean inequality [29] (AM-GM inequality):
$\forall a,b>0,\quad\alpha a+(1-\alpha)b\geq a^{\alpha}b^{1-\alpha},$ (7)
with equality if and only if $a=b$. Let $a=Z(\theta_{0})$ and $b=Z(\theta_{1})$ in the inequality of Eq. 7, we recover the inequality of Eq. 5. ∎
See the Appendix for the weaker proposition of a log-convex function being convex using the assumption of second-order differentiability of convex functions.
@@
The converse of Proposition 3 is not necessarily true: Some convex functions are not log-convex, and thus the class of strictly log-convex functions is a proper subclass of strictly convex functions. For example, $\theta^{2}$ is convex but log-concave since $(\log\theta^{2})^{\prime\prime}=-\frac{2}{\theta^{2}}<0$ (Figure 1).
Remark 1.
Since $Z=\exp(F)$ is strictly convex (Proposition 3), $F$ is exponentially convex.
Definition 1.
The cumulant function $F$ and partition function $Z$ of a regular exponential family are both strictly convex and smooth functions inducing a pair of dually flat spaces with corresponding Bregman divergences [9] $B_{F}$ (i.e., $B_{\log Z}$) and $B_{Z}$ (i.e., $B_{\exp F}$):
@@
and a pair of families of skewed Jensen divergences $J_{F,\alpha}$ and $J_{Z,\alpha}$:
@@
For a strictly convex function $F(\theta)$, we define the symmetric Jensen divergence:
$J_{F}(\theta_{1},\theta_{2})=J_{F,\frac{1}{2}}(\theta_{1}:\theta_{2})=\frac{F(\theta_{1})+F(\theta_{2})}{2}-F\left(\frac{\theta_{1}+\theta_{2}}{2}\right).$
Let $\mathcal{B}_{\Theta}$ denote the set of real-valued strictly convex and differentiable functions defined on an open set $\Theta$, called Bregman generators. We may equivalently consider the set of strictly concave and differentiable functions $G(\theta)$, and let $F(\theta)=-G(\theta)$: See [40] (Equation 1).
Remark 2.
@@
$\Leftrightarrow$ $\forall\theta_{1}\not=\theta_{2},F(\theta_{1})>F(\theta_{2})+{\left\langle\theta_{1}-\theta_{2},\nabla F(\theta)\right\rangle},$
@@
$\Leftrightarrow$ $\forall\theta_{1}\not=\theta_{2},\log Z(\theta_{1})>\log Z(\theta_{2})+{\left\langle\theta_{1}-\theta_{2},\frac{\nabla Z(\theta_{2})}{Z(\theta_{2})}\right\rangle}.$
@@
$J_{F,\alpha}^{s}(\theta_{1}:\theta_{2})=\left\{\begin{array}[]{ll}\frac{1}{\alpha(1-\alpha)}J_{F,\alpha}(\theta_{1}:\theta_{2}),&\alpha\in\mathbb{R}\backslash\{0,1\},\\ B_{F}(\theta_{1}:\theta_{2}),&\alpha=0,\\ 4\,J_{F}(\theta_{1},\theta_{2}),&\alpha=\frac{1}{2},\\ B_{F}^{*}(\theta_{1}:\theta_{2})=B_{F}(\theta_{2}:\theta_{1}),&\alpha=1.\end{array}\right.,$ (12)
@@
Remark 3.
Alternatively, we may rescale $J_{F}$ by a factor $\kappa(\alpha)=\frac{1}{\alpha(1-\alpha)4^{4\alpha(1-\alpha)}}$, i.e., $J_{F,\alpha}^{\bar{s}}(\theta_{1}:\theta_{2})=\kappa(\alpha)\,J_{F,\alpha}(\theta_{1}:\theta_{2})$ so that $\kappa(\frac{1}{2})=1$ and $J_{F,\frac{1}{2}}^{\bar{s}}(\theta_{1}:\theta_{2})=J_{F}(\theta_{1},\theta_{2})$.
Next, we first recall the connections between these Jensen and Bregman divergences which are divergences between parameters and statistical divergence counterparts between probability densities in §3. We will then introduce the novel connections between these parameter divergences and $\alpha$-divergences between unnormalized densities in §4.
@@ 3 Divergences related to the cumulant function
$D_{B,\alpha}^{s}(p:q)=-\frac{1}{\alpha(1-\alpha)}\log\int p^{\alpha}q^{1-\alpha}\mathrm{d}\mu,\quad\alpha\in\mathbb{R}\backslash\{0,1\}.$
The scaled $\alpha$-skewed Bhattacharyya distances can also be interpreted as Rényi divergences [38] scaled by $\frac{1}{\alpha}$: $D_{B,\alpha}^{s}(p:q)=\frac{1}{\alpha}D_{R,\alpha}(p:q)$ where
$D_{R,\alpha}(p:q)=\frac{1}{\alpha-1}\log\int p^{\alpha}q^{1-\alpha}\mathrm{d}\mu.$
The Bhattacharyya distance $D_{B}(p,q)=-\log\int\sqrt{pq}\mathrm{d}\mu$ corresponds to one fourth of $D_{B,\frac{1}{2}}^{s}(p:q)$: $D_{B}(p,q)=\frac{1}{4}\,D_{B,\frac{1}{2}}^{s}(p:q)$. Since $D_{B,\alpha}^{s}$ tend to the Kullback-Leibler divergence $D_{\mathrm{KL}}$ when $\alpha\rightarrow 1$ and to the reverse Kullback-Leibler divergence ${D_{\mathrm{KL}}}^{*}$ when $\alpha\rightarrow 0$, we have
$D_{B,\alpha}^{s}(p:q)=\left\{\begin{array}[]{ll}-\frac{1}{\alpha(1-\alpha)}\,\log\int p^{\alpha}q^{1-\alpha}\mathrm{d}\mu,&\alpha\in\mathbb{R}\backslash\{0,1\},\\ D_{\mathrm{KL}}(p:q),&\alpha=1,\\ 4\,D_{B}(p,q)&\alpha=\frac{1}{2},\\ {D_{\mathrm{KL}}}^{*}(p:q)=D_{\mathrm{KL}}(q:p)&\alpha=0.\end{array}\right.$
When both probability densities belong to the same exponential family =$\mathcal{E}=\{p_{\theta}(x)\ :\ \theta\in\Theta\}$ with cumulant $F(\theta)$, we have the following proposition:
Proposition 4 ([32]).
The scaled $\alpha$-skewed Bhattacharyya distances between two probability densities $p_{\theta_{1}}$ and $p_{\theta_{2}}$ of an exponential family amounts to the scaled $\alpha$-skewed Jensen divergence between their natural parameters:
@@
Proof.
The proof follows by first considering the $\alpha$-skewed Bhattacharyya similarity coefficient $\rho_{\alpha}(p,q)=\int p^{\alpha}q^{1-\alpha}\mathrm{d}\mu$.
$\rho_{\alpha}(p_{\theta_{1}}:p_{\theta_{2}})$ $=$ $\int\exp\left({\left\langle\theta_{1},x\right\rangle}-F(\theta_{1})\right)^{\alpha}\,\exp\left({\left\langle\theta_{2},x\right\rangle}-F(\theta_{2})\right)^{1-\alpha}\mathrm{d}\mu,$
$=$ $\int\exp({\left\langle\alpha\theta_{1}+(1-\alpha)\theta_{2}),x\right\rangle})\,\exp\left(-(\alpha F(\theta_{1})+(1-\alpha)F(\theta_{2}))\right)\mathrm{d}\mu.$
@@
Since $\bar{\theta}\in\Theta$, we have $\int\exp({\left\langle\bar{\theta},x\right\rangle}-F(\bar{\theta}))\mathrm{d}\mu=1$, and therefore $\rho_{\alpha}(p_{\theta_{1}}:p_{\theta_{2}})=\exp(-J_{F,\alpha}(\theta_{1}:\theta_{2}))$. ∎
@@
This formula has been further generalized in [31] by considering truncations of exponential family densities: Let $\mathcal{X}_{1}\subseteq\mathcal{X}_{2}\subseteq\mathcal{X}$ and $\mathcal{E}_{1}=\{1_{\mathcal{X}_{1}}(x)\,p_{\theta}(x)\}$, $\mathcal{E}_{2}=\{1_{\mathcal{X}_{2}}(x)\,q_{\theta^{\prime}}(x)\}$ be two truncated families of $\mathcal{X}$ with cumulant functions
$F_{1}(\theta)=\log\int_{\mathcal{X}_{1}}\exp({\left\langle t(x),\theta\right\rangle})\mathrm{d}\mu,$
and
$F_{2}(\theta^{\prime})=\log\int_{\mathcal{X}_{2}}\exp({\left\langle t(x),\theta\right\rangle})\mathrm{d}\mu\geq F_{1}(\theta^{\prime}).$
Then we have
@@
Truncated exponential families are normalized exponential families which may not be regular [14] (i.e., parameter space $\Theta$ may not be open).